Grade 119 grade maths

A+C=2B , then cosC-cosA / sinA-sinC should be equal to ?

Helper
8 Years agoGrade 11

1 Answer

Rajeswari
8 Years ago
A+c=2b ,cosc-cosa=-2sin(c+a)/2sin(a-c)/2÷2cos(c+a)/2sin(a-c)/2. =-2sin2b/2÷2cos2b/2. =sinb/cosb. Tanb